The most common isotope of hydrogen contains a proton and an electron separated by about 5.0 x 10-11 m. The mass of a proton is approximately 1.7 x 10-27 kg. The mass of the electron is approximately 9.0 x 10-31 kg.


1
Expert's answer
2022-02-17T10:05:21-0500

Given:

"m_1=1.7*10^{-27}\:\rm kg"

"m_2=9.0*10^{-31}\:\rm kg"

"r=5.0*10^{-11}\:\rm m"


The law of universal gravitation says


"F=G\frac{m_1m_2}{r^2}""F=6.67*10^{-11}\frac{1.7*10^{-27}*9.0*10^{-31}}{(5.0*10^{-11})^2}=4.1*10^{-47} \:\rm N"